OpenStudy (anonymous):

solve 5/12 m + 2/3 = -3/4

OpenStudy (anonymous):

\[\large \frac{5}{12}m+\frac{2}{3}=-\frac{3}{4}\]??

OpenStudy (anonymous):

yeah, sorry.

OpenStudy (anonymous):

no problem get rid of annoying fractions by multiplying both sides by \(12\)